OpenStudy (anonymous):

calculate dy/dx. You need not expand your answer. y = (9x^2 + x)(x − x^2)

OpenStudy (anonymous):

i used the product rule, its jsut getting all over the place

OpenStudy (anonymous):

by product rule, dy/dx = d(9x^2 + x)(x − x^2)+ (9x^2 + x)d(x − x^2) = (18x+1)(x-x^2) + (9x^2 +x)(1-2x)

OpenStudy (anonymous):

\[(9x^2 + x)(x-2x)+(18x+1)(x-x^2)\]

OpenStudy (paxpolaris):

and u can leave it at that

OpenStudy (anonymous):

so \[9x^3-18x^3+x^2-2x^2+18x^2-18x^3+1x-x^2\]?

OpenStudy (paxpolaris):

****need not expand your answer????

OpenStudy (anonymous):

Exapnd is to get \[(9x^2 + x)(x − x^2)=-9 x^4+8 x^3+x^2 \] and differentiate after.

OpenStudy (anonymous):

umm, im confused as to what you just said

OpenStudy (paxpolaris):

it should be \[\large(9x^2 + x)(1-2x)+(18x+1)(x-x^2)\]

OpenStudy (paxpolaris):

you wrote x-2x instead

OpenStudy (anonymous):

i understand, i am just having trouble solving that whole equation

OpenStudy (anonymous):

so ill try yours out right now

OpenStudy (anonymous):

so is it \[26x^2 - 48x^3 +2x \]

OpenStudy (anonymous):

which is \[-48x^3 + 26x^2 + 2x\] ???

OpenStudy (paxpolaris):

you don't need to ....the question says no need to expand... but if you do you should get\(-36x^3+24x^2+2x\)

OpenStudy (anonymous):

got it

OpenStudy (paxpolaris):

which is the same as if you differentiate \(y=-9 x^4+8 x^3+x^2\) as eliassaab suggested

OpenStudy (anonymous):

Without expansion \[(18 x+1) \left(x-x^2\right)+(1-2 x) \left(9 x^2+x\right) \]